OpenStudy (anonymous):

If the radius of a planet (not earth) is 3535 km, at what height above the surface of the planet will an astronaut lose 30% of his weight?

OpenStudy (anonymous):

W = GmM/r^2 W1/ W2 = (r2/r1)^2 = 0.70 solve for r1, given r2 = 3535 m

OpenStudy (anonymous):

Thank you!